# Ultra-Short Pulse Biphoton Source in Lithium Niobate Nanophotonics at 2 µm

### v1

@@ 3 Experimental Setup
The experimental setup used to characterize our biphoton source is shown in Fig.3. A 75-fs 250-MHz mode-locked laser centered at 1.045 µm is directed through a motorized half-wave plate (HWP) and polarizing beam splitter (PBS) combination used to digitally control the input power to the experiment. Laser light is then collected and focused onto the input facet of the waveguide via a reflective objective to minimize dispersion. Inside the waveguide, photons from the pump are split into signal/idler pairs via degenerate type-0 SPDC. Light from the output facet is collected via a lensed PM-2000 fiber from OZ Optics. This fiber is directed into freespace with a reflective coupler and passed through one 1330-nm long pass filter, three 1500-nm long pass filters, and a 48-nm bandpass filter centered at 2.09 µm to provide 150 dB of pump rejection. The photon pairs are then coupled back into fiber and passed through a series of fiber-coupled 50/50 splitters to perform Hanbury Brown-Twiss (HBT) and heralded $g_{H}^{(2)}(0)$ experiments for measuring the CAR, pair generation rate, and single photon behavior.
Fig. 4: A) Example histogram from an HBT experiment taken at 11 µW of on-chip pump power. The main coincidence peak, caused by a single entangled pair generation event, is centered at 0-ns delay while accidentals peaks, caused by uncorrelated pair generations, are present to the left and right at $\pm$4 ns. B) Count rates for the signal and idler channels compared with the fitted linear model.
Fig. 5: A) Coincidence-to-accidentals ratio (CAR) as a function of on-chip pump power B) On-chip pair generation rate as a function of on-chip pump power.
@@ 4 Measurement and Data Processing
To assess the CAR and on-chip pair generation rate, an HBT experiment (Fig.3B) is performed by passing photon pairs from the chip into a 50/50 beamsplitter and looking for coincidences at the output ports using a pair of superconducting nanowire single photon detectors (SNSPDs) from IDQuantique. True coincidences are created by entangled pairs of photons hitting the detectors at the same time, whereas accidentals are created by two photons from unrelated SNSPD processes causing coincidences. This is a measurement of the signal-to-noise ratio (SNR) of the detection system as the ratio of true to accidental coincidences is determined by the losses in the signal path, the dark count rates for the SNSPDs, and multi-photon generation events. Fig.4A shows a coincidence histogram collected by the time-to-digital converter recording events from the SNSPDs. The central peak at 0 ns is a result of both true and accidental coincidences, while side peaks are caused by accidentals. Hence, the CAR is defined as
@@
where $R_{si}$ is the total coincidence peak cont rate and $R_{acc}$ is the accidental peak count rate. Error bars are calculated via the standard deviation of CAR calculated from different accidentals peaks. The pair generation rate and detection efficiency for the signal and idler paths can be calculated using the simplified model:
@@
where $R_{s,i,si}$ are the signal, idler, and coincidence count rates respectively. $P$ is the power in mW, $\epsilon$ is the generation rate in pairs per mW, $\eta_{s,i}$ are the losses for the signal and idler paths including detector efficiency, and $R_{d_{s,i}}$ is the dark count for the signal and idler detectors. The factor of $\frac{1}{2}$ in Eq.4 is the result of a lack of deterministic signal/idler separation in our experiment as they are degenerate in wavelength and polarization. This model is valid for small values of $P$ where the pair generation rate is directly proportional to the input power for SPDC. In order to compute on-chip power for rate normalization, we measure the input loss by subtracting measurements of the output and throughput losses. The output loss is calculated by fitting the parametric generation output power vs pump power as detailed in the supplementary. Using this method, we calculate an output coupling loss of 10 dB from the chip to lensed fiber, 5 dB from the freespace filter, 3 dB from the transition from SM-2000 to SMF-28 fiber, and 6 dB from the SNSPD's detection efficiency for a total of 24 dB. This agrees well with the measured system efficiency of 25 $\pm$ 2 dB from fitting equations 2, 3, and 4.
Fig.5A displays the measured CAR values vs the on-chip pump power. We measure a maximum CAR of 945 $\pm$ 475 at a pump power of 100 nW. The large variance is a result of the lack of prominent accidental peaks at low pump powers. Fig.5B displays the measured on-chip pair generation rate as a function of the on-chip power. We measure a slope of 8.8 $\pm$ 2.3 GHz/mW over the entire SPDC spectrum. Using our filter bandwidth, this normalizes to a pair-generation rate of 440 $\pm$ 115 kHz/mW/GHz.
Fig. 6: Heralded $g_{H}^{(2)}(0)$ of our source as a function of on-chip pump power.
@@
where $R_{H,S,I}$ are the 2 and 3-fold coincidence rates of the heralding, signal, and idler detectors respectively. $P_{H,S,I}$ represent the probability for any given laser pulse to produce a click at the subscripted detectors [66]. Fig.6 shows the measured $g_{H}^{(2)}(0)$ as a function of the on-chip pump power. As the pump power decreases, $g_{H}^{(2)}(0)$ initially drops as fewer and fewer multiphoton events are observed. At very low pump powers, SNR is lost due to low count rates, causing $g_{H}^{(2)}(0)$ to rise again. We report a minimum $g_{H}^{(2)}(0)$ of 0.027$\pm$ 0.0084 at a pump power of 350 nW on chip, showing that we can reliably herald single photons from our device using the measurement setup. This performance is competitive with state-of-the-art mid and near-IR sources (see table 1 in the supplementary).
Fig. 7: A comparison plot of relevant photon pair sources. Horizontal error bars represent the reported source bandwidth. Data for efficiency, wavelength, and bandwidth are taken from [16, 30, 18, 67] are PPLN ridge waveguides, similar to our device. [17, 68, 69] are larger (>1-µm tall) PPLN waveguides. [33] is a bulk PPLN crystal. [36] is a bulk AlGaS2 crystal. [32] is a nanophotonic AlN ring resonator. [70] is a bulk LBO crystal. [71] is an InGaP ring resonator.
@@ 5 Comparison and Discussion
Fig.7 compares a variety of different IR pair sources with our work. We include 1550-nm sources to demonstrate that not only does our device have state-of-the-art performance at 2 µm, but it is also competitive with state-of-the-art devices for standard telecommunications bands. This also allows for a more fair comparison of pair generation rate as most reported 2-µm sources utilize the $\chi^{(3)}$ instead of the $\chi^{(2)}$ nonlinearity, which has a normalized pair generation rate relative to $mW^{2}$ of pump power instead of $mW$. Table 1 in the supplementary goes into greater detail by normalizing the efficiency by the device length and bandwidth for an apples-to-apples comparison. Looking at the last pair generation rate column ($Hz/mW/cm^{2}/GHz$), we see that our generation rate measurements agree well with other devices reported in literature except for one recent demonstration which exceeds the normalized pair generation of all other listed sources by an order of magnitude[18]. It should be noted that the pair generation rate will intrinsically be lower at 2 µm compared to 1550 nm as a result of the $\lambda^{-4}$ dependence [39]. In the lower portion of the table which lists 2-µm sources, we see that our source exceeds the CAR, pair generation rate, and heralded $g_{H}^{(2)}(0)$ of the other recent 2-µm demonstrations in bulk crystals. We attribute this namely to the use of nanophotonic TFLN, which is responsible for the stronger nonlinear interactions given the device length, mode confinement, and dispersion engineering.
@@ 6 Conclusion
We have demonstrated a dispersion engineered broadband biphoton source with a record breaking CAR, pair generation rate, and heralded $g_{H}^{(2)}(0)$ in the 2-µm window. The near-zero GVM and GVD operation of our source allows for use of ultrashort pulses, a crucial resource for realizing large-scale quantum information processing systems as the temporal confinement of these pulses can be used to both enhance the strength of nonlinear interactions as well as create well-defined time bins for temporal multiplexing of quantum gates beyond 1-THz clock speeds. Reduced solar irradiance near 2-µm combined with technological advancements in transmission and detection in the mid-IR make our source an ideal candidate for both freespace and fiber based quantum key distribution. Future work will revolve around characterizing the quantum interference, temporal width, and entanglement properties of the source through Houng-Ou-Mandel (HOM) and Franson interferometry. We plan to improve outcoupling losses via both on-chip adiabatic couplers for pump rejection and the use of an inverse tapered waveguide for better mode matching with a lensed fiber. Using in-house adiabatic coupler designs and state-of-the-art inverse tapers [72], we can lower the output coupling losses to 0.32 dB and filter losses to 3 dB for a total detection loss of 9.32 dB.
Device nanofabrication was performed at the Kavli Nanoscience Institute (KNI) at Caltech.
The authors gratefully acknowledge support from ARO grant no. W911NF-23-1-0048, NSF grant no. 1846273 and 1918549, AFOSR award FA9550-20-1-0040, the center for sensing to intelligence at Caltech, and NASA/JPL. The authors wish to thank NTT Research for their financial support.
All authors have accepted responsibility for the entire content of this manuscript and approved its submission.
L.L., R.S and A.M. are involved in developing photonic integrated nonlinear circuits at PINC Technologies Inc. L.L. and A.M. have an equity interest in PINC Technologies Inc. The other authors declare that they have no competing interests.
Informed consent was obtained from all individuals included in this study.
The conducted research is not related to either human or animals use.
Data sets used in this paper are available from the corresponding authors upon request.

### v2

@@ 4 Experimental Setup
The experimental setup used to characterize our biphoton source is shown in Fig.3. A 75-fs 250-MHz mode-locked laser centered at 1.045 µm is directed through a motorized half-wave plate (HWP) and polarizing beam splitter (PBS) combination used to digitally control the input power to the experiment. Laser light is then collected and focused onto the input facet of the waveguide via a reflective objective to minimize dispersion. Inside the waveguide, photons from the pump are split into signal/idler pairs via degenerate type-0 SPDC. Light from the output facet is collected with a lensed PM-2000 fiber from OZ Optics. This fiber is directed into freespace with a reflective coupler and passed through one 1330-nm long pass filter, three 1500-nm long pass filters, and a 48-nm bandpass filter centered at 2.09 µm to provide 150 dB of pump rejection. The photon pairs are then coupled back into fiber and passed through a series of fiber-coupled 50/50 splitters to perform Hanbury Brown-Twiss (HBT) and heralded $g_{H}^{(2)}(0)$ experiments for measuring the CAR, pair generation rate, and single photon behavior.
Figure 4: A) Example histogram from an HBT experiment taken at 11 µW of on-chip pump power. The main coincidence peak, caused by the detection of a signal/idler pair from the same generation event, is centered at 0-ns delay while accidentals peaks, caused by signal and idler photons created from consecutive and hence uncorrelated generation events, are present to the left and right at $\pm$4 ns. B) Count rates for the signal and idler channels compared with the fitted linear model.
Figure 5: A) Coincidence-to-accidentals ratio (CAR) as a function of on-chip pump power. The fitted model is taken from [prabhakar2020two]. B) On-chip pair generation rate as a function of on-chip pump power. This is the rate of pair generation at the output of the poled region of the waveguide. Pump input coupling losses are calculated based on a parametric gain measurement detailed in the supplementary.
@@ 5 Measurement and Data Processing
To assess the CAR and on-chip pair generation rate, an HBT experiment (Fig.3B) is performed by passing photon pairs from the chip into a 50/50 beamsplitter and looking for coincidences at the output ports using a pair of superconducting nanowire single photon detectors (SNSPDs) from IDQuantique. True coincidences are created by entangled pairs of photons hitting the detectors at the same time, whereas accidentals are created by two photons from unrelated SNSPD processes causing coincidences. This is a measurement of the signal-to-noise ratio (SNR) of the detection system as the ratio of true to accidental coincidences is determined by the losses in the signal path, the dark count rates for the SNSPDs, and multi-photon generation events. Fig.4A shows a coincidence histogram collected by the time-to-digital converter recording events from the SNSPDs. The central peak at 0 ns is a result of both true and accidental coincidences, while side peaks are caused by accidentals. For a continuous-wave pump, accidental counts are spread evenly across the delay histogram. For the pulsed case, accidentals counts are localized to multiples of the pump repetition time (4 ns in our case for our 250-MHz repetition rate). The CAR is defined as
@@
where $R_{si}$ is the total coincidence peak cont rate and $R_{acc}$ is the accidental peak count rate. Error bars are calculated via the standard deviation of CAR calculated from different accidentals peaks. The timing jitter in the SNSPD measurement is limited by the jitter of the electronics (100 ps), resulting in coincidence and accidentals peaks which are much wider than the temporal width of the biphotons themselves. Fig.4B shows the count rates at the signal and idler detectors as a function of the on-chip pump power. By combining individual detector count rates with the coincidence count rates, the pair generation rate on-chip and detection efficiency for the signal and idler paths can be calculated from fitting the simplified linear model:
@@
where $R_{s,i,si}$ are the signal, idler, and coincidence count rates respectively. $P$ is the pump power in mW, $\epsilon$ is the generation rate in pairs per mW, $\eta_{s,i}$ are the losses for the signal and idler paths including detector efficiency, and $R_{d_{s,i}}$ is the dark count for the signal and idler detectors. The factor of $\frac{1}{2}$ in Eq.4 is the result of a lack of deterministic signal/idler separation in our experiment as they are degenerate in wavelength and polarization. This model is valid for small values of $P$ where the pair generation rate is directly proportional to the input power for SPDC. To compute on-chip power for rate normalization, we measure the input loss by subtracting measurements of the output lensed fiber and throughput losses. The output loss for the lensed fiber is calculated by fitting the parametric generation output power vs pump power as detailed in the supplementary. Using this method, we calculate an output coupling loss of 9 dB from the chip to lensed fiber. The losses for additional components are measured with a 2 µm diode laser. We measure 5 dB from the freespace filter, 3 dB from the transition from SM-2000 to SMF-28 fiber, and 6 dB from the SNSPD’s detection efficiency for a total of 23 dB. This agrees well with the measured total system efficiency of 25 $\pm$ 2 dB from fitting experimental data to equations 2, 3, and 4.
Fig.5A displays the measured CAR values vs the on-chip pump power which has been fitted with a model provided by [prabhakar2020two]. At high pump powers, the CAR is limited by noise induced from multiphoton events generated by a strong pump pulse. As the pump power is lowered, the CAR increases until the maximum SNR is achieved. The CAR then decreases at lower pump powers due to a loss of signal altogether from the reduced pair generation rate. We measure a maximum CAR of 945 $\pm$ 475 at a pump power of 100 nW. The large variance is a result of the lack of prominent accidental peaks at low pump powers. The theoretical model used to fit the CAR vs power curve depends only on the pair generation rate, detector dark count rates, and system losses. A perfectly achromatic detection system would measure a higher CAR in the absence of the filter due to the higher observed pair generation rate. This is not the case for our system as our SNSPDs and fibers suffer higher losses at longer wavelengths, and therefore a bandpass filter in the measurement path increases our measured CAR by limiting the spectral dependence of the measurement efficiency to help satisfy the assumptions of the model. Fig.5B displays the measured on-chip pair generation rate as a function of the on-chip power. This is the rate at which generated photon pairs exit the poled region of the waveguide before incurring detection losses. We measure a slope of 8.8 $\pm$ 2.3 GHz/mW over the entire SPDC spectrum. Using our filter bandwidth, this normalizes to a pair-generation rate of 440 $\pm$ 115 kHz/mW/GHz.
Figure 6: Heralded $g_{H}^{(2)}(0)$ of our source as a function of on-chip pump power.
@@
where $R_{H,S,I}$ are the 2 and 3-fold coincidence rates of the heralding, signal, and idler detectors respectively. $P_{H,S,I}$ represent the probability for any given laser pulse to produce a click at the subscripted detectors [signorini2020chip]. Fig.6 shows the measured $g_{H}^{(2)}(0)$ as a function of the on-chip pump power. As the pump power decreases, $g_{H}^{(2)}(0)$ initially drops as fewer and fewer multiphoton events are observed. At very low pump powers, SNR is lost due to low count rates, causing $g_{H}^{(2)}(0)$ to rise again. We report a minimum $g_{H}^{(2)}(0)$ of 0.027$\pm$ 0.0084 at a pump power of 3.5 µW on chip, showing that we can reliably herald single photons from our device using the measurement setup. This performance is competitive with state-of-the-art mid and near-IR sources (see table 1 in the supplementary). The CAR and $g_{H}^{(2)}(0)$ performance are maximized at two different pump powers because the $g_{H}^{(2)}(0)$ measurement uses 3 detectors (as opposed to 2 for CAR) and is therefore more susceptible to system losses, necessitating a higher power to maximize the system SNR. Increasing overall system losses would also shift the maximum CAR to a higher power as these losses translate to a lower overall signal.
Figure 7: A comparison plot of relevant photon pair sources. Horizontal error bars represent the reported source bandwidth. Data for efficiency, wavelength, and bandwidth are taken from [javid2021ultrabroadband, zhao2020high, xue2021ultrabright, bock2016highly] are PPLN ridge waveguides, similar to our device. [zhang2023scalable, yoshizawa2003generation, mori2006distribution] are larger (¿1-µm tall) PPLN waveguides. [prabhakar2020two] is a bulk PPLN crystal. [kumar2021mid] is a bulk AlGaS2 crystal. [guo2017parametric] is a nanophotonic AlN ring resonator. [noh2006noncollinear] is a bulk LBO crystal. [zhao2022ingap] is an InGaP ring resonator.
@@ 6 Comparison and Discussion
Fig.7 compares a variety of different IR pair sources with our work. We include 1550-nm sources to demonstrate that not only does our device have state-of-the-art performance at 2 µm, but it is also competitive with state-of-the-art devices for standard telecommunications bands. This allows for a more fair comparison of pair generation rate as most reported 2-µm sources utilize the $\chi^{(3)}$ instead of the $\chi^{(2)}$ nonlinearity, which has a normalized pair generation rate relative to $mW^{2}$ of pump power instead of $mW$. Table 1 in the supplementary goes into greater detail by normalizing the efficiency with the device length and bandwidth for an apples-to-apples comparison. Looking at the last pair generation rate column ($Hz/mW/cm^{2}/GHz$), we see that our rate measurements agree well with other devices reported in literature except for one recent demonstration which exceeds the normalized pair generation of all other listed sources by an order of magnitude[xue2021ultrabright]. It should be noted that the pair generation rate will intrinsically be lower at 2 µm compared to 1550 nm as a result of the $\lambda^{-4}$ dependence [jankowski2021dispersion]. In the lower portion of the table which lists 2-µm sources, we see that our source exceeds the CAR, pair generation rate, and heralded $g_{H}^{(2)}(0)$ of the other recent 2-µm demonstrations in bulk crystals. We attribute this namely to the use of nanophotonic TFLN, which is responsible for the stronger nonlinear interactions given the device length, mode confinement, and dispersion engineering.
@@ 7 Conclusion
We have demonstrated a dispersion engineered broadband biphoton source with a record breaking CAR, pair generation rate, and heralded $g_{H}^{(2)}(0)$ in the 2-µm window. The near-zero GVM and GVD operation of our source allows for use of ultrashort pulses, a crucial resource for realizing large-scale quantum information processing systems as the temporal confinement of these pulses can be used to both enhance the strength of nonlinear interactions as well as create well-defined time bins for temporal multiplexing of quantum gates beyond 1-THz clock speeds. Reduced solar irradiance near 2-µm combined with technological advancements in transmission and detection in the mid-IR make our source an ideal candidate for both freespace and fiber based quantum key distribution. Future work will revolve around characterizing the quantum interference, temporal width, and entanglement properties of the source through Houng-Ou-Mandel (HOM) and Franson interferometry. We plan to improve outcoupling losses via both on-chip adiabatic couplers for pump rejection and the use of an inverse tapered waveguide for better mode matching with a lensed fiber. Using in-house adiabatic coupler designs and state-of-the-art inverse tapers [yao2020efficient], we can lower the output coupling losses to 0.32 dB and filter losses to 3 dB for a total detection loss of 9.32 dB.
8 Acknowledgements and Disclosures
Device nanofabrication was performed at the Kavli Nanoscience Institute (KNI) at Caltech. The authors gratefully acknowledge support from ARO grant no. W911NF-23-1-0048, NSF grant no. 1846273 and 1918549, AFOSR award FA9550-20-1-0040, the center for sensing to intelligence at Caltech, and NASA/JPL. The authors wish to thank NTT Research for their financial support. All authors have accepted responsibility for the entire content of this manuscript and approved its submission. L.L. and A.M. are involved in developing photonic integrated nonlinear circuits at PINC Technologies Inc. L.L. and A.M. have an equity interest in PINC Technologies Inc. The other authors declare that they have no competing interests. Informed consent was obtained from all individuals included in this study. The conducted research is not related to either human or animals use. Data sets used in this paper are available from the corresponding authors upon request.
[…]